# NumGrad-Pull: Numerical Gradient Guided Tri-plane Representation for Surface Reconstruction from Point Clouds

### v1

@@ Abstract
Reconstructing continuous surfaces from unoriented and unordered 3D points is a fundamental challenge in computer vision and graphics. Recent advancements address this problem by training neural signed distance functions to pull 3D location queries to their closest points on a surface, following the predicted signed distances and the analytical gradients computed by the network. In this paper, we introduce NumGrad-Pull, leveraging the representation capability of tri-plane structures to accelerate the learning of signed distance functions and enhance the fidelity of local details in surface reconstruction. To further improve the training stability of grid-based tri-planes, we propose to exploit numerical gradients, replacing conventional analytical computations. Additionally, we present a progressive plane expansion strategy to facilitate faster signed distance function convergence and design a data sampling strategy to mitigate reconstruction artifacts. Our extensive experiments across a variety of benchmarks demonstrate the effectiveness and robustness of our approach. Code is available at https://github.com/CuiRuikai/NumGrad-Pull.
1 Introduction
@@
The signed distance function (SDF) has become a fundamental 3D representation in computer vision and graphics, widely adopted for its utility in various applications. Recent advances [27, 22, 8, 2, 40] have increasingly focused on leveraging SDFs to reconstruct continuous surfaces from 3D point clouds [9, 35, 36, 10]. By mapping each 3D coordinate to a corresponding signed distance from the shape surface, SDFs implicitly define surfaces as zero-level sets, providing a powerful surface representation framework that offers distinct advantages in capturing high-fidelity 3D shapes with complex, arbitrary topologies.
Recent efforts [27, 2, 20, 22, 19, 28] in learning SDFs from point clouds tend to directly train neural networks on point sets, relying on extensive distance computations between query and surface points to guide the network towards accurate distance field convergence. A pioneering approach in this direction is Neural-Pull [27], which trains a neural network to iteratively project query points onto surfaces, using the learned signed distances and gradients to guide each time of adjustment. Subsequent works have extended this method by incorporating additional constraints for improved pulling target estimation [22, 20, 2] or enhanced scalability of the approach [8]. In general, existing methods often utilize deeper neural architectures and complex loss constraints to model signed distance fields, yet they still produce overly smoothed surfaces lacking fine-grained details and suffer from slow query speeds.
@@
2 Related Work
2.1 Traditional Surface Reconstruction
2.2 Neural Surface Reconstruction
@@
More recently, neural networks have enabled learning-based approaches to surface reconstruction, where implicit functions trained on large datasets facilitate detailed surface modeling even with sparse inputs [29, 32]. However, estimating surfaces from unoriented point sets remains challenging without ground truth signed distance values. Point2Surf [13] addresses this issue by learning a prior over local patches with detailed structures and coarse global information. Moreover, Ma et al. [27] introduce the first pulling operation-based surface reconstruction scheme, observing that the gradient of a neural SDF represents the normal direction of a query point. By using this gradient and the estimated distance from the SDF, any query point can be pulled to the surface, considering the neural SDF as a close approximation of the ground truth. Building on this scheme, Grid-Pull [8] presents a discrete grid-based representation to improve the scalability of pull-based methods. Subsequent work by Koneputugodage et al. [20] applies guiding points for incrementally changing optimizations toward the true surfaces. More recently, Li et al. [22] propose a bilateral filter to smooth the implicit field, retaining high-frequency geometric details. Different from existing studies, we are pioneering the exploration of hybrid explicit–implicit representations [5] for fundamental surface reconstruction problems, where our approach shows promise in accelerating query speeds and preserving geometric details across various reconstruction tasks [24, 12].
2.3 Neural surface representation
@@
Implicit representations have become widely used for surface reconstruction from point clouds, e.g., Neural-Pull [27] devises a deep fully-connected network to represent the distance field as a continuous function. However, this approach results in slow query speeds, as each distance query requires a complete pass through the network. In contrast, explicit representations, such as discrete voxel grids [8], offer faster query times but face memory limitations, with memory consumption scaling as $O(n^{3})$ as the discretization resolution increases. Given complementary advantages provided by explicit and implicit representations, motivated to develop a hybrid approach for surface reconstruction that combines the strengths of both paradigms. In this work, we present the first hybrid explicit–implicit method for neural surface reconstruction from point clouds, leveraging the tri-plane structure [5, 31, 30, 23, 12, 11, 39] to integrate the efficiency of explicit storage with the flexibility of implicit decoding. Particularly, our proposed tri-plane method stores features on axis-aligned planes and exploys a lightweight implicit decoder to aggregate these features for efficient querying, achieving computational and memory efficiency.
3 Method
@@
NumGrad-Pull reconstructs continuous surfaces from unoriented point sets by parameterizing a signed distance function (SDF) with a hybrid explicit-implicit tri-plane structure (Sec. 3.2). The network is trained with numerical gradients (Sec. 3.3), which help stabilize the training process. Additionally, we introduce a progressive tri-plane expansion strategy (Sec. 3.4) and a novel query location sampling method (Sec. 3.5) to enhance both convergence speed and reconstruction quality. The overall pipeline of our proposed method is illustrated in Fig. 2.
3.1 Preliminaries
@@
A signed distance function is a function $\Phi:\mathbb{R}^{3}\mapsto\mathbb{R}$ that maps a spatial position to a real value representing the signed distance to the shape surface. Using this function, the shape surface can be defined as the zero-level set of $\Phi$, allowing for surface extraction through methods such as marching cubes [26]. To learn a neural SDF $\Phi$ from a given point cloud $P$ without normal information, Ma et al. [27] propose to query the function at various positions, and train a network to “pull” each query position $q$ toward the shape surface, approximated by the nearest point on $P$ to $q$. Specifically, this pulling operation is defined as:
$q^{\prime}=q-|\Phi(q)|\times\frac{\nabla\Phi(q)}{\|\nabla\Phi(q)\|},$ (1)
@@ 3.3 Numerical Gradient
After performing bilinear interpolation on each of the three planes ($XY$, $XZ$, and $YZ$), the final feature vector is obtained by aggregating the features from each plane:
@@
[…]
@@
3.4 Progressive Tri-Plane Expansion
@@
3.5 Query Locations Sampling
@@
However, this data sampling strategy is insufficient for the learning of our tri-plane-based hybrid SDF representations. The abovementioned procedure guides the object’s surface, while neglecting distant regions. Since our method encodes spatial information on a tri-plane structure, this lack of sufficient guidance can lead to inaccurate or random predictions for the signed distances in distant areas. To address this issue, we introduce another complementary sampling strategy of randomly sampling points in a unit cube $[-1,1]^{3}$, aligning with the scale of training data. Accordingly, we not only ensure that every tri-plane entity is trained, but also prevent untrained regions of the tri-plane grid from negatively affecting the learning process.
3.6 Optimization
@@ 4 Experiments
4.1 Implementation Details
We initialize the tri-plane with a resolution of $8^{2}$ and a feature dimension of 32. The tri-planes are then expanded over three stages at the 3k, 8k, and 12k iterations, ultimately reaching a resolution of $32^{2}$. To decode the tri-plane features, we use a three-layer fully connected network with a hidden dimension of 128. Additionally, we initialize the network parameters using the geometric network initialization scheme proposed in Gropp et al. [14], which approximates the signed distance function of a sphere. For optimization, we use the Adam optimizer [17], with an initial learning rate of 0.001 for the MLP and 0.05 for the tri-plane parameters. All the experiments are conducted on an Nvidia GeForce RTX 3090 GPU.
Table 1: Comparisons on ABC [18] and FAMOUS [13] datasets. The metric is chamfer distance ($CD_{\ell_{2}}$) scaled by $10^{3}$.
@@ 4.1 Implementation Details
Table 2: Comparisons on the ShapeNet dataset [6]. The metric is chamfer distance ($CD_{\ell_{2}}$) scaled by $10^{3}$.
Method Venue Display Lamp Plane Cabinet Vessel Table Chair Sofa Avg.
SPSR [16] SGP 2006 2.730 2.270 2.170 3.630 2.540 3.830 2.930 2.760 2.860
NP [27] ICML 2021 0.390 0.800 0.080 0.260 0.220 0.600 0.540 0.120 0.380
LPI [7] ECCV 2022 0.080 0.172 0.060 0.179 0.092 0.436 0.187 0.164 0.171
PCP [28] CVPR 2022 0.887 0.380 0.065 0.153 0.079 0.131 0.110 0.086 0.136
GP [8] ICCV 2023 0.082 0.347 0.007 0.112 0.033 0.052 0.043 0.015 0.086
IF [22] ECCV 2024 0.009 0.019 0.045 0.055 0.005 0.025 0.070 0.027 0.032
Ours N/A 0.024 0.013 0.012 0.032 0.012 0.022 0.024 0.020 0.020
Table 3: Comparison on the Surface Reconstruction Benchmark [38].
Method Anchor Daratech DC Gargoyle Lord Quas $CD_{\ell_{1}}$ $HD$ $d_{\overrightarrow{C}}$ $d_{\overrightarrow{H}}$ $CD_{\ell_{1}}$ $HD$ $d_{\overrightarrow{C}}$ $d_{\overrightarrow{H}}$ $CD_{\ell_{1}}$ $HD$ $d_{\overrightarrow{C}}$ $d_{\overrightarrow{H}}$ $CD_{\ell_{1}}$ $HD$ $d_{\overrightarrow{C}}$ $d_{\overrightarrow{H}}$ $CD_{\ell_{1}}$ $HD$ $d_{\overrightarrow{C}}$ $d_{\overrightarrow{H}}$ SPSR [16] 0.60 14.89 0.60 14.89 0.44 7.24 0.44 7.24 0.27 3.10 0.27 3.10 0.26 6.80 0.26 6.80 0.20 4.61 0.20 4.61 IGR [14] 0.22 4.71 0.12 1.32 0.25 4.01 0.08 1.59 0.17 2.22 0.09 2.61 0.16 3.52 0.06 0.81 0.12 1.17 0.07 0.98 SIREN [37] 0.32 8.19 0.10 2.43 0.21 4.30 0.09 1.77 0.15 2.18 0.06 2.76 0.17 4.64 0.08 0.91 0.17 0.82 0.12 0.76 VisCo [34] 0.21 3.00 0.15 1.07 0.21 4.06 0.14 1.76 0.15 2.22 0.09 2.76 0.17 4.40 0.11 0.96 0.12 1.06 0.07 0.64 SAP [33] 0.12 2.38 0.08 0.83 0.26 0.87 0.04 0.41 0.07 1.17 0.04 0.53 0.07 1.49 0.05 0.78 0.05 0.98 0.04 0.51 NP [27] 0.122 3.243 0.061 3.208 0.375 3.127 0.746 3.267 0.157 3.541 0.242 3.523 0.080 1.376 0.063 0.475 0.064 0.822 0.053 0.508 GP [8] 0.093 1.804 0.066 0.460 0.062 0.648 0.039 0.293 0.066 1.103 0.036 0.539 0.063 1.129 0.045 0.700 0.047 0.569 0.031 0.370 DIGS [2] 0.063 1.447 0.030 0.270 0.049 0.858 0.025 0.441 0.042 0.667 0.022 0.729 0.047 0.971 0.028 0.271 0.031 0.496 0.017 0.181 IF [22] 0.052 1.232 0.025 0.265 0.051 0.751 0.028 0.423 0.041 0.815 0.019 0.724 0.044 1.089 0.022 0.246 0.030 0.554 0.014 0.230 Ours 0.051 1.194 0.041 0.113 0.050 0.730 0.036 0.236 0.026 0.215 0.032 0.286 0.033 0.416 0.040 0.127 0.036 0.293 0.039 0.127
Table 4: Effect of different tri-plane configurations on surface reconstruction for the FAMOUS dataset [13] compared with recent methods, measured by $CD_{\ell_{2}}$ and scaled by $10^{4}$.
Metric 8 16 32 64 NP [27] IF [22] $CD_{\ell_{2}}$downarrow 9.31 0.45 0.39 0.35 1.00 0.80 Speed (iter/s) $\uparrow$ 219 231 227 225 122 42
@@
4.2 Surface Reconstruction for Shapes
4.2.1 Datasets and Metrics
For surface reconstruction of general shapes from raw point clouds, we follow the evaluation protocol used in prior work [27, 22] and assess our method on three widely used datasets: ShapeNet [6], ABC [18], and FAMOUS [13]. For ShapeNet, we use the test split defined in [25] and the pre-processed data provided in [29]. Following the footsteps of Erler et al. [13], we employ the same train/test split and pre-processed data for the ABC and FAMOUS datasets.
@@ 4.2.1 Datasets and Metrics
4.2.2 Comparison
We compare the performance of our proposed NumGrad-Pull with several state-of-the-art methods, including SPSR [16], PCP [28], LPI [7], SIREN [37], Neural-Pull [27], DIGS [2], Grid-Pull [8], and the recent IF [22]. Results on the ABC and FAMOUS datasets are presented in Tab. 1, highlighting the superiority of our method: NumGrad-Pull outperforms the latest state-of-the-art by margins of 0.02 and 0.04 on the ABC and FAMOUS datasets, respectively. Furthermore, our method shows strong performance in Tab. 2 on the ShapeNet dataset, which is a challenging and comprehensive benchmark comprising over 3,000 testing objects. Our NumGrad-Pull achieves the best average surface reconstruction quality across all methods, outperforming the state-of-the-art IF [22] method by 0.012 and achieving the best results in six out of eight categories. These quantitative results demonstrate the effectiveness of our approach in improving surface reconstruction quality.
@@ 4.2.2 Comparison
4.3 Surface Reconstruction for Real Scans
4.3.1 Datasets and Metrics
@@
4.3.2 Comparison
We compare our method with various prior methods using the specified evaluation metrics, with per-object results presented in Tab. 3. Our method performs well across multiple metrics and objects, particularly achieving superior reconstruction quality on the Anchor and the Gargoyle object. To further substantiate the results, we provide visualizations of the Gargoyle (7th column) and DC (8th column) samples in Fig. 3. Our method captures finer details, such as the rings on the Gargoyle’s wings, and preserves accurate geometry even in areas with corruptions in the original scan, demonstrating robust detail retention and reconstruction fidelity.
4.4 Ablation Studies

### v2

@@ Abstract
Reconstructing continuous surfaces from unoriented and unordered 3D points is a fundamental challenge in computer vision and graphics. Recent advancements address this problem by training neural signed distance functions to pull 3D location queries to their closest points on a surface, following the predicted signed distances and the analytical gradients computed by the network. In this paper, we introduce NumGrad-Pull, leveraging the representation capability of tri-plane structures to accelerate the learning of signed distance functions and enhance the fidelity of local details in surface reconstruction. To further improve the training stability of grid-based tri-planes, we propose to exploit numerical gradients, replacing conventional analytical computations. Additionally, we present a progressive plane expansion strategy to facilitate faster signed distance function convergence and design a data sampling strategy to mitigate reconstruction artifacts. Our extensive experiments across a variety of benchmarks demonstrate the effectiveness and robustness of our approach.
I Introduction
The signed distance function (SDF) has become a fundamental 3D representation for visual and spatial data modalities, widely adopted in various multimedia applications such as shape modeling, animation, virtual/augmented reality, etc. Recent advances [24, 19, 8, 2] have increasingly focused on leveraging SDFs to reconstruct continuous surfaces from 3D point clouds. By mapping each 3D coordinate to a corresponding signed distance from the shape surface, SDFs implicitly define surfaces as zero-level sets, providing a powerful surface representation framework that offers distinct advantages in capturing high-fidelity 3D shapes with complex, arbitrary topologies.
Recent efforts [24, 2, 16, 19, 17, 25, 36] in learning SDFs from point clouds tend to directly train neural networks on point sets, relying on extensive distance computations between query and surface points to guide the network towards accurate distance field convergence. A pioneering approach in this direction is Neural-Pull [24], which trains a neural network to iteratively project query points onto surfaces, using the learned signed distances and gradients to guide each time of adjustment. Subsequent works have extended this method by incorporating additional constraints for improved pulling target estimation [19, 16, 2] or enhanced scalability of the approach [8]. In general, existing methods often utilize deeper neural architectures and complex loss constraints to model signed distance fields, yet they still produce overly smoothed surfaces lacking fine-grained details and suffer from slow query speeds.
@@ II-A Traditional Surface Reconstruction
II-B Neural Surface Reconstruction
@@
More recently, neural networks have enabled learning-based approaches to surface reconstruction, where implicit functions trained on large datasets facilitate detailed surface modeling even with sparse inputs [26, 29]. However, estimating surfaces from unoriented point sets remains challenging without ground truth signed distance values. Point2Surf [10] addresses this issue by learning a prior over local patches with detailed structures and coarse global information. Moreover, [24] introduce the first pulling operation-based surface reconstruction scheme, observing that the gradient of a neural SDF represents the normal direction of a query point. By using this gradient and the estimated distance from the SDF, any query point can be pulled to the surface, considering the neural SDF as a close approximation of the ground truth. Building on this scheme, Grid-Pull [8] presents a discrete grid-based representation to improve the scalability of pull-based methods. Subsequent work by [16] applies guiding points for incrementally changing optimizations toward the true surfaces. More recently, [19] propose a bilateral filter to smooth the implicit field, retaining high-frequency geometric details. Different from existing studies, we are pioneering the exploration of hybrid explicit–implicit representations [5] for fundamental surface reconstruction problems, where our approach shows promise in accelerating query speeds and preserving geometric details across various reconstruction tasks [21, 9, 38, 37, 32].
II-C Neural surface representation
@@
Implicit representations have become widely used for surface reconstruction from point clouds, e.g., Neural-Pull [24] devises a deep fully-connected network to represent the distance field as a continuous function. However, this approach results in slow query speeds, as each distance query requires a complete pass through the network. In contrast, explicit representations, such as discrete voxel grids [8], offer faster query times but face memory limitations, with memory consumption scaling as $O(n^{3})$ as the discretization resolution increases. Given complementary advantages provided by explicit and implicit representations, motivated to develop a hybrid approach for surface reconstruction that combines the strengths of both paradigms. In this work, we present the first hybrid explicit–implicit method for neural surface reconstruction from point clouds, leveraging the tri-plane structure [5, 28, 27, 20] to integrate the efficiency of explicit storage with the flexibility of implicit decoding. Particularly, our proposed tri-plane method stores features on axis-aligned planes and deploys a lightweight implicit decoder to aggregate these features for efficient querying, achieving computational and memory efficiency.
III Method
@@
NumGrad-Pull reconstructs continuous surfaces from unoriented point sets by parameterizing a signed distance function (SDF) with a hybrid explicit-implicit tri-plane structure (Sec. III-B). The network is trained with numerical gradients (Sec. III-C), which help stabilize the training process. Additionally, we introduce a progressive tri-plane expansion strategy (Sec. III-D) and a novel query location sampling method (Sec. III-E) to enhance both convergence speed and reconstruction quality. The overall pipeline of our proposed method is illustrated in Fig. 2.
III-A Preliminaries
@@
A signed distance function is a function $\Phi:\mathbb{R}^{3}\mapsto\mathbb{R}$ that maps a spatial position to a real value representing the signed distance to the shape surface. Using this function, the shape surface can be defined as the zero-level set of $\Phi$, allowing for surface extraction through methods such as Marching Cubes [23]. To learn a neural SDF $\Phi$ from a given point cloud $P$ without normal information, [24] propose to query the function at various positions, and train a network to “pull” each query position $q$ toward the shape surface, approximated by the nearest point on $P$ to $q$. Specifically, this pulling operation is defined as:
$q^{\prime}=q-\Phi(q)\times\frac{\nabla\Phi(q)}{\|\nabla\Phi(q)\|},$ (1)
@@ III-C Numerical Gradient
After performing bilinear interpolation on each of the three planes ($XY$, XZXZ, and $YZ$), the final feature vector is obtained by aggregating the features from each plane:
@@
However, the feature extraction in the forward pass involves non-differentiable rounding operations, such as $\lfloor\cdot\rfloor$ and $\lceil\cdot\rceil$. As a result, the gradient of the tri-plane encoding is inherently local: when the query point $q$ crosses grid boundaries, the corresponding grid corner entities change discontinuously. Consequently, the supervision loss is only back-propagated to the grid entities involved in the current forward pass, i.e., the four corner points involved in the interpolation. This local gradient propagation becomes problematic when learning neural SDFs from sparse point clouds. In such cases, the sparsity of the points can confuse the network about surface connectivity as the discrete grid features do not imply continuous gradient changes, leading to corrupted learning.
@@
III-D Progressive Tri-Plane Expansion
@@
III-E Query Locations Sampling
@@
However, this data sampling strategy is insufficient for the learning of our tri-plane-based hybrid SDF representations. The abovementioned procedure guides the object’s surface, while neglecting distant regions. Since our method encodes spatial information on a tri-plane structure, this lack of sufficient guidance can lead to inaccurate or random predictions for the signed distances in distant areas. To address this issue, we introduce another complementary sampling strategy of randomly sampling points in a unit cube $1,1]^{3}$, aligning with the scale of training data. Accordingly, we not only ensure that every tri-plane entity is trained, but also prevent untrained regions of the tri-plane grid from negatively affecting the learning process.
III-F Optimization
@@ IV Experiments
IV-A Implementation Details
We initialize the tri-plane with a resolution of $8^{2}$ and a feature dimension of 32. The tri-planes are then expanded over three stages at the 3k, 8k, and 12k iterations, ultimately reaching a resolution of $32^{2}$. To decode the tri-plane features, we use a three-layer fully connected network with a hidden dimension of 128. Additionally, we initialize the network parameters using the geometric network initialization scheme proposed in [11], which approximates the signed distance function of a sphere. For optimization, we use the Adam optimizer [14], with an initial learning rate of 0.001 for the MLP and 0.05 for the tri-plane parameters. All the experiments are conducted on an Nvidia GeForce RTX 3090 GPU.
TABLE I: Comparisons on ABC [15] and FAMOUS [10] datasets. The metric is chamfer distance ($CD_{\ell_{2}}$) scaled by $10^{3}$.
@@ IV-A Implementation Details
TABLE II: Comparisons on the ShapeNet dataset [6]. The metric is chamfer distance ($CD_{\ell_{2}}$) scaled by $10^{3}$.
[…]
TABLE III: Comparison on the Surface Reconstruction Benchmark [34].
[…]
@@
Figure 4: Visual comparisons of surface reconstruction quality on the 3D Scene dataset [35].
IV-B Surface Reconstruction for Shapes
IV-B1 Datasets and Metrics
For surface reconstruction of general shapes from raw point clouds, we follow the evaluation protocol used in prior work [24, 19] and assess our method on three widely used datasets: ShapeNet [6], ABC [15], and FAMOUS [10]. For ShapeNet, we use the test split defined in [22] and the pre-processed data provided in [26]. Following the footsteps of [10], we employ the same train/test split and pre-processed data for the ABC and FAMOUS datasets.
@@ IV-B1 Datasets and Metrics
IV-B2 Comparison
We compare the performance of our proposed NumGrad-Pull with several state-of-the-art methods, including SPSR [13], PCP [25], LPI [7], SIREN [33], Neural-Pull [24], DIGS [2], Grid-Pull [8], and the recent IF [19]. Results on the ABC and FAMOUS datasets are presented in Tab. I, highlighting the superiority of our method: NumGrad-Pull outperforms the latest state-of-the-art by margins of 0.02 and 0.04 on the ABC and FAMOUS datasets, respectively. Furthermore, our method shows strong performance in Tab. II on the ShapeNet dataset, which is a challenging and comprehensive benchmark comprising over 3,000 testing objects. Our NumGrad-Pull achieves the best average surface reconstruction quality across all methods, outperforming the state-of-the-art IF [19] method by 0.012 and achieving the best results in six out of eight categories. These quantitative results demonstrate the effectiveness of our approach in improving surface reconstruction quality.
@@ IV-B2 Comparison
IV-C Surface Reconstruction for Real Scans
IV-C1 Datasets and Metrics
@@
IV-C2 Comparison
We compare our method with various prior methods using the specified evaluation metrics, with per-object results presented in Tab. III. Our method performs well across multiple metrics and objects, particularly achieving superior reconstruction quality on the Anchor and the Gargoyle object. To further substantiate the results, we provide visualizations of the Gargoyle (7th column) and DC (8th column) samples in Fig. 3. Our method captures finer details, such as the rings on the Gargoyle’s wings, and preserves accurate geometry even in areas with corruptions in the original scan, demonstrating robust detail retention and reconstruction fidelity.
IV-D Qualitative Results on 3D Scenes
We further evaluate surface reconstruction on the real-world 3D Scene dataset [35]. Fig. 4 presents qualitative comparisons on three representative scenes. Compared with NeuralPull [24], GridPull [8], and IF [19], our approach consistently produces higher-quality surfaces. Moreover, fine structures like thin parts and sharp boundaries are better preserved, leading to clearer separation among nearby instances in crowded scenes.
This improvement can be attributed to our tri-plane design. Prior MLP-only SDFs are primarily supervised using nearest-point target pairs, where the supervision signal is highly local and lacks global constraints, particularly for large, scene-level scans. In our NumGrad-Pull method, we leverage dense near-surface Gaussian samples to learn accurate local surface details, and additionally exploits uniformly sampled points within the unit cube to provide global free-space supervision for tri-plane learning. As a result, our tri-plane design not only encodes local geometry from dense near-surface supervision, but also leverages uniform samples for global regularization, leading to more stable, robust, and generalizable reconstruction performance on real-world 3D scenes.
IV-E Ablation Studies